OpenStudy (anonymous):

Rosa is setting up a model train track that is 5 5/6 feet long. Along the track, she places a telephone pole every 5/12 foot apart. How many telephone poles does she need? 7 poles 35 poles 12 poles 14 poles

OpenStudy (ganpat):

5 5/6 is nothing but.. 35/6.. and one pole at a distance of 5/12.. so (35/6) / (5/12) = 35 / 6 * 12 / 5 = 14, i guess.. i just calculated orally, u recheck again..